# A Scalable Measure of Loss Landscape Curvature for Analyzing the Training Dynamics of LLMs

### v1

@@ 4 How much Pre-training data is needed to avoid Catastrophic forgetting?
To go beyond demonstrating pre-existing sharpness phenomena at scale, we next provide a new practical application of critical sharpness, focused on data mixing. We are motivated by the fact that during finetuning, neural networks are prone to catastrophic forgetting, where the model performance degrades on the pretraining dataset and benchmarks as the model adapts to the new task (McCloskey and Cohen, 1989; Luo et al., 2025; McLeish et al., 2025). To mitigate this, several strategies have been proposed (De Lange et al., 2022), with mixing samples from the pre-training data being the most effective (Robins, 1995). This practice is reflected in large-scale training: for example, Walsh et al. (2025) use a mid-training mix consisting of approximately $50\%$ pre-training data. Intuitively, we want to use as much as possible new data in fine-tuning, using as little pre-training rehearsal data as possible to retain base capabilities. However, it remains unclear what fraction of pre-training data is sufficient to effectively prevent catastrophic forgetting. To this end, we leverage critical sharpness to systematically examine the effect of adding pre-training data to the training mix during OLMo mid-training and provide actionable guidance for selecting the pre-training fraction without exhaustive grid search.
The goal of mid-training or finetuning is to improve the performance on specialized domains or tasks while preserving generic capabilities acquired during pre-training. Intuitively, this requires the model to adapt to the new task, while staying within the “pre-training basin” —a region of the parameter space where the pre-training loss remains low. Leaving this basin would be marked by an increase in the pre-training loss, and the critical learning rate quantifies exactly how far we can step before this occurs. To formalize this intuition, we define relative critical learning rate and sharpness, as follows:
@@
Consider a model with parameters $\bm{\theta}$, two loss functions $L_{1}(\bm{\theta})$ and $L_{2}(\bm{\theta})$, and an update direction $\Delta\bm{\theta}$ derived from $L_{2}$. The relative critical learning rate is the smallest learning rate for which taking a step in the direction $\Delta\bm{\theta}_{2}$ increases the loss $L_{1}(\bm{\theta})$:
$\eta_{c}^{1\to 2}=\min_{\eta>0}\{\eta\mid L_{1}(\bm{\theta}-\eta\Delta\bm{\theta}_{2})>L_{1}(\bm{\theta})\}.$ (6)
@@
This definition is general and applies whenever optimization on one objective may affect performance on another. The two losses can correspond to different tasks (e.g., pre-training vs. finetuning), different loss functions (e.g., next-token prediction vs. reinforcement learning objectives), or different data characteristics (e.g., short vs. long context). In this work, we focus on the finetuning setting, where $L_{1}({\bm{\theta}})$ is the pre-training loss on a general text corpus and $L_{2}({\bm{\theta}})$ is the finetuning loss on a specialized dataset such as math.
@@
To assess the impact of pre-training data fraction in the training mix during fine-training, we consider the OLMo-2 77B pre-trained checkpoint as our starting point. We compute the update direction from a mixture composed of DCLM (pre-training data) and the math subset of Dolmino mix (Walsh et al., 2025), and examine how the relative critical sharpness varies with the DCLM fraction in this mixture [^3] When measuring the relative critical sharpness, we do not update the model parameters..
@@
Figure˜6(a) shows how the relative critical sharpness varies with the DCLM ratio in the training mix for several evaluation tasks. When the mix contains mostly math data (low DCLM ratio), the sharpness for DCLM is an order of magnitude higher than for other tasks, indicating that the pre-training loss landscape is particularly sharp and sensitive to finetuning updates in this regime. As more pre-training data is added, the relative critical sharpness for most tasks decreases, suggesting that the landscapes align. On the other hand, when the DCLM ratio approaches one, the sharpness for downstream tasks such as Math, GSM8K, and MMLU increases, meaning that these tasks become the limiting factor for the maximum stable learning rate. Notably, there is an intermediate DCLM ratio (around 0.7) where the sharpness curves for different tasks intersect, representing a sweet spot that allows for the largest possible learning rate without being constrained by any single task.
@@
Figure 6: (a) Relative critical sharpness (Definition˜4.1) for various evaluation tasks. The shaded region around the mean trends denotes the variation across batches. (b, c) GSM8K and MMLU accuracy as a function of pre-training (DCLM) mix ratio and learning rate. Red indicates a decrease in performance relative to the checkpoint, white indicates no change, and blue indicates an improvement. The black dashed line denotes the smallest critical learning rate among at the pre-trained checkpoint across tasks from (a) and white dashed line denotes the critical learning rate for the corresponding task.
@@
In Figure˜6(b, c), we evaluate the impact of the pre-training mix by training the pre-trained checkpoint for 11B tokens and measuring downstream accuracy on GSM8K and MMLU benchmarks. In these plots, red indicates a decrease in performance relative to the checkpoint, white indicates no change, and blue indicates an improvement. We observe a natural trade-off: GSM8K accuracy is typically maximized outside the pre-training basin (i.e., with low DCLM ratio and high learning rate), but this comes at the expense of MMLU performance, which is best preserved within the pre-training basin. Therefore, if the sole objective is to improve on the finetuning task (math), then training primarily on it with a large learning rate is effective, though it may lead to forgetting on other tasks (MMLU). In contrast, if the goal is to improve on the finetuning task (math) while maintaining pre-training performance (MMLU), it is essential to include pre-training data in the mix, with a sweet spot emerging around a DCLM ratio of $0.6$ and a learning rate of 3e3e-$05$ in our experiment. Remarkably, this is close to the optimum of $0.7$ suggested by the critical sharpness analysis. This result highlights the importance of balancing task-specific and general data in the training mix. Once the pre-trained loss $L_{1}$ increases ($\eta>2/\lambda_{c}^{1\to 2}$), there is no guarantee that further updates will improve $L_{1}$ again, unless the pre-train data is injected into the mix.
@@ 5 Related Works
In neural network training, if the learning rate exceeds the threshold $\sim 2/\lambda^{H}_{\max}$, it causes the loss to increase in the next training step (Wu et al., 2018). However, Kalra and Barkeshli (2023) demonstrated that the empirical critical learning rate can be much higher than this theoretical threshold derived from convex analysis, with critical learning rate reaching up to $40/\lambda^{H}_{\max}$ when sharpness decreases significantly during training. Our work extends theirs by showing that a similarly large critical learning rate can also arise when sharpness increases during training, due to contributions from other eigendirections (Figure˜3). Beyond its role in characterizing training instabilities, Hessian sharpness also exhibits progressive sharpening during training. In particular, Cohen et al. (2021) showed that sharpness increases when the learning rate is decayed, which we corroborate with additional analysis at large scales (Figures˜4 and 5).
@@
The works most closely related to ours are those by Kalra and Barkeshli (2024) and Roulet et al. (2024). Kalra and Barkeshli (2024) used critical learning rate to set the initial learning rate during warmup. By comparison, Roulet et al. (2024) proposed setting the learning rate at ‘edge’ throughout training, i.e., $\eta_{t}=2/\lambda_{\text{dir}}$, and show that it can outperform constant learning rate training, but not typical learning rate schedules, consisting of learning rate warmup followed by decay. Vaswani et al. (2019) also study setting the learning rate as $\eta=1/\lambda_{\text{dir}}$. In contrast, we use the critical learning rate to study the curvature dynamics of large-scale models and examine the effect of pre-training data on catastrophic forgetting.
@@
Catastrophic forgetting is a long-standing problem in neural networks, first identified by McCloskey and Cohen (1989), where model performance degrades on previously learned tasks as it adapts to the new task. Approaches to mitigate catastrophic forgetting can be categorized into (i) regularization methods (Kirkpatrick et al., 2017; Ahn et al., 2019), (ii) ensembling and parameter isolation (Rusu et al., 2016; Aljundi et al., 2017), and (iii) rehearsal (Robins, 1995; Lopez-Paz and Ranzato, 2017). Among these, rehearsal —mixing samples from the pre-training data —has become the most widely adopted strategy due to its simplicity and effectiveness. Recent works have demonstrated that catastrophic forgetting persists in LLM finetuning (Luo et al., 2025; Huang et al., 2024; Scialom et al., 2022). Particularly related to our work is that of Chen et al. (2025), who show that if finetuning retains previously learned capabilities, the model remains in “most-case” and “worst-case” basins.
@@ 6 Discussion and Conclusion
We also introduced relative critical sharpness, which quantifies the curvature of one loss landscape along the update direction of another. Using this measure, we identified a sweet spot in the pre-training data fraction that balances specialization and retention during finetuning, enabling practitioners to evaluate data composition choices without extensive ablations. Beyond finetuning, relative critical sharpness provides a general framework for analyzing changes in the loss landscape due to distribution shifts, changes in loss functions, or modifications to the training data mixture.

### v2

@@ 4 How much Pre-training data is needed to avoid Catastrophic Forgetting?
To go beyond demonstrating pre-existing sharpness phenomena at scale, we next provide a new practical application of critical sharpness, focused on data mixing. We are motivated by the fact that during fine-tuning, neural networks are prone to catastrophic forgetting, where the model performance degrades on the pretraining dataset and benchmarks as the model adapts to the new task (MCCLOSKEY1989109; lou2023_forgetting; mcleish2025teachingpretrainedlanguagemodels). To mitigate this, several strategies have been proposed (Lange2022), with mixing samples from the pre-training data being the most effective (ROBINS01061995). This practice is reflected in large-scale training: for example, walsh2025 uses a mid-training mix consisting of approximately $50\%$ pre-training data. Intuitively, we want to use as much new data as possible in fine-tuning, while using as little pre-training rehearsal data as possible to retain base capabilities. However, it remains unclear what fraction of pre-training data is sufficient to effectively prevent catastrophic forgetting. To this end, we leverage critical sharpness to systematically examine the effect of adding pre-training data to the training mix during OLMo mid-training and provide actionable guidance for selecting the pre-training data fraction without exhaustive grid search.
The goal of mid-training or fine-tuning is to improve the performance on specialized domains or tasks while preserving generic capabilities acquired during pre-training. Intuitively, this requires the model to adapt to the new task, while staying within the “pre-training basin” —a region of the parameter space where the pre-training loss remains low. Leaving this basin would be marked by an increase in the pre-training loss, and the critical learning rate quantifies exactly how far we can step before this occurs. To formalize this intuition, we define relative critical learning rate and sharpness, as follows:
@@
Consider a model with parameters $\bm{\theta}$, two loss functions $L_{1}(\bm{\theta})$ and $L_{2}(\bm{\theta})$, and an update direction $\Delta\bm{\theta}$ derived from $L_{2}$. The relative critical learning rate is the smallest learning rate for which taking a step in the direction $\Delta\bm{\theta}$ increases the loss $L_{1}(\bm{\theta})$:
$\eta_{c}^{1\to 2}=\inf\{\eta>0\mid L_{1}(\bm{\theta}-\eta\Delta\bm{\theta})>L_{1}(\bm{\theta})\}.$ (6)
@@
This definition is general and applies whenever optimization on one objective may affect performance on another. The two losses can correspond to different tasks (e.g., pre-training vs. fine-tuning), different loss functions (e.g., next-token prediction vs. reinforcement learning objectives), or different data characteristics (e.g., short vs. long context). In this work, we focus on the fine-tuning setting, where $L_{1}({\bm{\theta}})$ is the pre-training loss on a general text corpus and $L_{2}({\bm{\theta}})$ is the fine-tuning loss, which often is a mixture of pre-training and specialized data (e.g., math).
@@
To assess the impact of pre-training data fraction in the training mix during fine-tuning, we consider the OLMo-2 77B pre-trained checkpoint as our starting point. We define $L_{2}(\bm{\theta})$ as the loss on a mixture composed of DCLM (pre-training data) and the math subset of Dolmino mix (walsh2025), and compute the update direction $\Delta\bm{\theta}$ from this mixture. We then examine how the relative critical sharpness varies with the DCLM fraction in this mixture [^3] When measuring the relative critical sharpness, we do not update the model parameters..
@@
Figure˜6(a) shows how the relative critical sharpness varies with the DCLM ratio in the training mix for several evaluation tasks. When the mix contains mostly math data (low DCLM ratio), the sharpness for DCLM is an order of magnitude higher than for other tasks, indicating that the pre-training loss landscape is particularly sharp and sensitive to fine-tuning updates in this regime. As more pre-training data is added, the relative critical sharpness for most tasks decreases, suggesting that the landscapes align. On the other hand, when the DCLM ratio approaches one, the sharpness for downstream tasks such as Math, GSM8K, and MMLU increases, meaning that these tasks become the limiting factor for the maximum stable learning rate. Notably, there is an intermediate DCLM ratio (around 0.7) where the sharpness curves for different tasks intersect, representing a sweet spot that allows for the largest possible learning rate without being constrained by any single task.
@@
Figure 6: (a) Relative critical sharpness (Definition˜4.1) for various evaluation tasks ($L_{1}$). The shaded region around the mean trends denotes the variation across batches. (b, c) GSM8K and MMLU accuracy as a function of pre-training (DCLM) mix ratio and learning rate. Red indicates a decrease in performance relative to the checkpoint, white indicates no change, and blue indicates an improvement. The black dashed line denotes the smallest critical learning rate at the pre-trained checkpoint across tasks from (a), and the white dashed line denotes the critical learning rate for the corresponding task.
@@
In Figure˜6(b, c), we evaluate the impact of the pre-training mix by training the pre-trained checkpoint for 11B tokens and measuring downstream accuracy on GSM8K and MMLU benchmarks. In these plots, red indicates a decrease in performance relative to the pre-trained checkpoint, white indicates no change, and blue indicates an improvement. We observe a natural trade-off: GSM8K accuracy is typically maximized outside the pre-training basin (i.e., with low DCLM ratio and high learning rate), but this comes at the expense of MMLU performance, which is best preserved within the pre-training basin. Therefore, if the sole objective is to improve on the fine-tuning task (math), then training primarily on it with a large learning rate is effective, though it may lead to forgetting on other tasks (MMLU). In contrast, if the goal is to improve on the fine-tuning task (math) while maintaining pre-training performance (MMLU), it is essential to include pre-training data in the mix, with a sweet spot emerging around a DCLM ratio of $0.6$ and a learning rate of 3e3e-$05$ in our experiment. Remarkably, this is close to the optimum of $0.7$ suggested by the critical sharpness analysis. This result highlights the importance of balancing task-specific and general data in the training mix. Once the pre-trained loss $L_{1}$ increases ($\eta>2/\lambda_{c}^{1\to 2}$), there is no guarantee that further updates will improve $L_{1}$ again, unless the pre-training data is injected into the mix.
@@ 5 Related Works
In neural network training, if the learning rate exceeds the threshold $\sim 2/\lambda^{H}_{\max}$, it causes the loss to increase in the next training step (Wu2018_how_sgd). However, kalra2023phase demonstrated that the empirical critical learning rate can be much higher than this theoretical threshold derived from convex analysis, with critical learning rate reaching up to $40/\lambda^{H}_{\max}$ when the sharpness naturally decreases during training. Our work extends theirs by showing that a similarly large critical learning rate can also arise when sharpness increases during training, due to contributions from other eigendirections (Figure˜3). Beyond its role in characterizing training instabilities, Hessian sharpness also exhibits progressive sharpening during training. In particular, cohen2021gradient showed that sharpness increases when the learning rate is decayed, which we corroborate with additional analysis at large scales (Figures˜4 and 5).
@@
The works most closely related to ours are those by kalra2024why, roulet2024stepping and bu2025gradient. kalra2024why used the critical learning rate to set the initial learning rate during warmup. By comparison, roulet2024stepping proposed setting the learning rate at ‘edge’ throughout training, i.e., $\eta_{t}=2/\lambda_{\text{dir}}$, and show that it can outperform constant learning rate training, but not typical learning rate schedules, consisting of learning rate warmup followed by decay. Vaswani2019Painless also study setting the learning rate as $\eta=1/\lambda_{\text{dir}}$. bu2025gradient proposed a generalized Newton method that uses forward passes to estimate local curvature for automatic learning rate selection. In contrast, we use the critical learning rate to study the curvature dynamics of large-scale models and examine the effect of pre-training data on catastrophic forgetting.
@@
Catastrophic forgetting is a long-standing problem in neural networks, first identified by MCCLOSKEY1989109, where model performance degrades on previously learned tasks as it adapts to the new task. Approaches to mitigate catastrophic forgetting can be categorized into (i) regularization methods (Kirkpatrick2017; Ahn2019), (ii) ensembling and parameter isolation (rusu2016progressive; Aljundi2017), and (iii) rehearsal (ROBINS01061995; Lopez-Paz2017). Among these, rehearsal —mixing samples from the pre-training data —has become the most widely adopted strategy due to its simplicity and effectiveness. Recent works have demonstrated that catastrophic forgetting persists in LLM fine-tuning (lou2023_forgetting; huang-etal-2024-mitigating; scialom-etal-2022-fine). Particularly related to our work is that of chen2025understandingpretrainingfinetuningloss, who show that if fine-tuning retains previously learned capabilities, the model remains in “most-case” and “worst-case” basins. Theoretical works have further elucidated these dynamics; tahir2025features argue that fine-tuning success relies on high feature overlap with the pre-trained model, while graldi2025the show that staying in the “lazy” training regime can reduce forgetting. In the light of these works, relative critical sharpness can be used to quantify the feature overlap between the pre-trained model and the fine-tuning dataset and the critical learning rate required to stay in the lazy regime.
@@ 6 Discussion and Conclusion
We also introduced relative critical sharpness, which quantifies the curvature of one loss landscape along the update direction of another. Using this measure, we identified a sweet spot in the pre-training data fraction that balances specialization and retention during fine-tuning, enabling practitioners to evaluate data composition choices without extensive ablations. Beyond fine-tuning, relative critical sharpness provides a general framework for analyzing changes in the loss landscape due to distribution shifts, changes in loss functions, or modifications to the training data mixture.